OpenStudy (anonymous):

what is the integral of 1/(xln(x))dx please show work

OpenStudy (lgbasallote):

Hint: let u = lnx

OpenStudy (anonymous):

ohhhhhhhh thank you i feel really dumb now

OpenStudy (lgbasallote):

hehe..it takes practice to see these kinds of things ^_^ just practice :)

OpenStudy (anonymous):

ya ill get it eventually I hope but hey so since u=lnx then du = 1/x so would it substitute to be 1/u du cause that doesn't seem right

OpenStudy (lgbasallote):

yup that's about right... du = 1/x dx so \(\int \frac{1}{x \ln x} dx = \int \frac{1}{u} du\)

OpenStudy (anonymous):

ok ya nevermind i'm sorry its kinda late where I am that is like calc one stuff but thank you for helping my confusion